OpenStudy (anonymous):

Find the derivative of y=arctan[radical(x^2 -1)] + arccsc(x) using implicit differentation. x>1

hartnn (hartnn):

hi :) lets first simplify the 1st term = arctan[radical(x^2 -1)] put x = tan u what u get ?

OpenStudy (anonymous):

so am i just finding the derivative of the first part? I got \[\frac{ 2x }{ (x ^{2}-1)^2 }\] Not sure if thats right; i used \[\frac{ d }{ dx }\tan ^{-1}u=\frac{ 1 }{ 1+u ^{2} }\times\frac{ d }{ dx }u\]

hartnn (hartnn):

you missed the radical part ? anyways, we can first simplify the 1st term before taking the derivative if you don't simplify, it'll be quite complicated

OpenStudy (anonymous):

oh ok, and my bad; i forgot about the radical haha does it come out to \[\frac{ 2x }{ x^{2} } = \frac{ 2 }{ x }\] ?

OpenStudy (anonymous):

i don't know how to find the derivative of arc csc. I only know how if its something like y=arc csc(x) csc(y)=x

OpenStudy (yttrium):

I think it is easy if you use quotient rule. Even without simplifying. Do you know how to apply it @claborte ?

OpenStudy (anonymous):

do i use quotient rule for the whole problem or just the arc csc part?

OpenStudy (yttrium):

Ow wait! I didn't see the true problem.

OpenStudy (anonymous):

yeah, its a doozy.

OpenStudy (raffle_snaffle):

The derivative of csc(x)=csc(x)cot(x)